# AgriFM: A Multi-source Temporal Remote Sensing Foundation Model for Crop Mapping

### v1

@@ Abstract
Escalating climate change and global population growth demand precise crop mapping solutions to ensure food security. Accurate crop mapping fundamentally relies on modeling multi-scale spatiotemporal patterns, where spatial scales range from individual field textures to landscape-level context, and temporal scales capture both short-term phenological transitions and full growing-season dynamics, demanding robust joint spatiotemporal modeling capabilities. Existing approaches are inherently limited by their architectural designs that process spatial and temporal features separately, leading to significant information loss. Transformer-based remote sensing foundation models (RSFMs) offer promising potential for crop mapping due to their innate ability for unified spatiotemporal processing via self-attention mechanisms. However, current RSFMs remain suboptimal for crop mapping: they either employ fixed spatiotemporal windows that ignore the multi-scale nature of crop systems or completely disregard temporal information by focusing solely on spatial patterns. To bridge these gaps, we present AgriFM, a multi-source remote sensing foundation model specifically designed for agricultural crop mapping. Our approach begins by establishing the necessity of simultaneous hierarchical spatiotemporal feature extraction, leading to the development of a modified Video Swin Transformer architecture where temporal down-sampling is synchronized with spatial scaling operations. This modified backbone enables efficient unified processing of long time-series satellite inputs while preserving critical multi-scale spatial patterns and phenological dynamics. AgriFM leverages temporally rich data streams from three satellite sources including MODIS, Landsat-8/9 and Sentinel-2, and is pre-trained on a global representative dataset comprising over 25 million image samples supervised by land cover products. The resulting framework incorporates a versatile decoder architecture that dynamically fuses these learned spatiotemporal representations, supporting diverse downstream tasks including cropland mapping, field-boundary delineation, early-season crop mapping, and specific crop mapping (e.g., winter wheat and paddy rice) with difference data sources. Comprehensive evaluations demonstrate AgriFM’s superior performance over conventional deep learning approaches and state-of-the-art general-purpose RSFMs across all downstream tasks. Notably, AgriFM demonstrates remarkable extensibility by supporting multi-source remote sensing data and variable temporal windows for diverse mapping tasks. Codes will be available at https://github.com/flyakon/AgriFM and https://glass.hku.hk.
@@ 1 Introduction
The growing demands for accurate crop mapping have spurred significant methodological evolution, beginning with phenology-based approaches as the foundational paradigm. They utilize the distinct physical attributes of various crops, including their differing reflectance across spectral bands and their unique phenological traits at various growth phases. Examining and detecting the alterations throughout the crop growth cycle can efficiently differentiate various crops (Qiu et al., 2022; Dong et al., 2020; Qiu et al., 2017; Liang et al., 2024; Ashourloo et al., 2022). Qiu et al. (Qiu et al., 2022) introduced an innovative approach for generating yearly 500-m MODIS-derived national maps of various cropping systems in China, employing phenology-based mapping algorithms and pixel purity thresholds, resulting in an overall accuracy of 89%. Ashourloo et al. (Ashourloo et al., 2022) introduced a novel phenology-based approach utilizing Sentinel-2 time-series data for the efficient differentiation of wheat and barley across extensive regions, attaining an overall accuracy exceeding 76%. The issue with these methods is that they require setting different thresholds based on the mapping task and study area, which significantly affects their applicability. For complex scenarios and tasks, this may lead to erroneous or even failed mapping results.
@@
Transformer-based RSFMs have demonstrated superior capabilities in extracting spatiotemporal features across various remote sensing tasks, significantly outperforming traditional machine learning methods and deep learning architectures. This suggests RSFMs hold substantial potential for crop mapping, as they can theoretically model complete crop growth cycles end-to-end while integrating multi-source remote sensing data to improve mapping accuracy. Several pioneering studies have begun exploring RSFMs for crop mapping tasks. Fang et al. (Fang et al., 2025) developed a rice mapping method using NASA-IBM’s Prithvi foundation model (Jakubik et al., 2023) with Harmonized Landsat and Sentinel-2 (HLS30) data (Claverie et al., 2018) and successfully generating rice distribution maps across Monsoon Asia. Similarly, Qi et al. (Qin et al., 2025) proposed a spatiotemporal masking strategy for pretraining a spatiotemporal collaborative learning network (STCLN) to extract informative representations for crop mapping.
@@
While these approaches demonstrate promising results, they predominantly rely on Vision Transformer (ViT) architectures. This preference stems from ViT’s straightforward temporal data processing and relatively lower computational demands during masked image modeling pretraining. However, ViT’s inherent downsampling mechanism and fixed spatiotemporal patch windows may adversely affect crop mapping performance. Previous research (Chen et al., 2017; Huang et al., 2020) has shown pixel-wise classification tasks require high spatial feature fidelity, yet ViT’s patch embedding operation typically employs aggressive and fixed downsampling (e.g., 16×16 patches), potentially compromising subtle spectral-spatial differences. Although Fang and Qi’s implementations incorporate specialized modifications to mitigate these limitations, such adaptations increase methodological complexity while limiting generalizability.
@@
In contrast, Swin Transformer’s hierarchical feature extraction capability - resembling CNN’s multi-scale processing - appears better suited for crop mapping tasks. Like U-Net architectures, Swin Transformer can effectively fuse features across scales to enhance spatial precision (Cao et al., 2022). However, existing Swin-based RSFMs remain scarce and typically ignore temporal information, severely limiting their applicability for crop monitoring. While Video Swin Transformer (Liu et al., 2022) inherently supports joint spatiotemporal processing, no existing RSFM has yet adopted it as the backbone architecture.
@@
To address these gaps, this paper first establishes the necessity of simultaneous hierarchical spatiotemporal feature extraction for crop mapping, leading to the development of a modified Video Swin Transformer (Liu et al., 2022) architecture where temporal down-sampling is synchronized with spatial scaling operations. This modified backbone enables efficient unified processing of long time-series satellite inputs while preserving critical multi-scale spatial patterns and phenological dynamics. We then develop AgriFM, a multi-source temporal remote sensing foundation model specifically designed for agricultural crop mapping.
@@
AgriFM leverages temporally rich data streams from MODIS, Landsat-8 / 9, and Sentinel-2 satellites, pre-trained on a global representative dataset with supervision from global land cover products. The pretraining dataset comprises over 25 million globally sampled images from MODIS (250m & 500m), Landsat-8/9 (30m), and Sentinel-2 (10m & 20m). For each satellite source, we randomly selected 16 chronological frames per year to capture complete seasonal variations. Unlike RSFMs relying solely on self-supervision, we incorporate land cover supervision using GLC_FCS30D data (Zhang et al., 2024c). Specifically, we extract image-level land cover fractions within each sample area as regression targets, optimizing with L1 loss. This approach follows established practices demonstrating land cover priors’ effectiveness (Li et al., 2022; Wang et al., 2024). To address inherent label noise in reference products, we implemented an auxiliary teacher network with exponential moving average updates, effectively guiding the model toward robust feature learning while filtering out unreliable supervision signals.
@@
AgriFM incorporates a versatile decoder that dynamically fuses extracted multi-scale spatiotemporal representations, supporting diverse crop mapping tasks, including cropland mapping, field-boundary delineation, early-season crop mapping and specific crop mapping (e.g., winter wheat and paddy rice). For systematic evaluation, we compare AgriFM with three representative approaches: (1) ViT-based temporal RSFMs (Prithvi (Jakubik et al., 2023), SatMAE (Cong et al., 2022)), (2) Swin-based models without temporal processing (PIS (An et al., 2024)), and (3) conventional deep learning methods (CNNs / LSTMs). Experimental results demonstrate AgriFM’s consistent superiority, particularly in two crucial aspects: preserving fine-grained spatial details essential for field-level mapping, and effectively modeling long-term temporal patterns critical for crop phenology analysis - capabilities where existing methods show notable limitations. AgriFM achieves unified processing of heterogeneous data sources, variable temporal sequences, and diverse mapping tasks within a single network architecture. This comprehensive method overcomes the traditional need for task-specific model adaptations while delivering robust performance across crop mapping tasks.
@@ 2.1.1 Multi-source Satellite Data for Pre-training
The MODIS data from the Terra satellite include the surface reflectance products MOD09A1 and MOD09Q1 with the temporal resolution of 8 days and the spatial resolutions of 500 m and 250 m, respectively. The MOD09Q1 product comprises the first two bands at 250m, and we extract bands 3-7 at 500m from the MOD09A1 product and then re-sample them to 250 meters. The MODIS data were selected for this process from 2020 and 2021.
@@ 2.1.3 Pre-training Dataset
After the pre-processing, we have constructed a multi-source temporal remote sensing pre-training dataset, which is derived from MODIS, Landsat 8/9, and Sentinel-2 satellite data at 250m&500m, 30m, and 10m&20m spatial resolution. To meet the input requirements of deep learning models, we randomly select these satellite data and crop the original satellite data into images of 224 x 224 size. For each image, a sequence of time series data is also generated with a minimum length of 16. Land cover fractions are computed to serve as pre-training supervision and the image with only background land cover is abandoned.
@@ 2.2 Method
In this section, we introduce our proposed multi-source temporal remote sensing foundation model for crop mapping. AgriFM is built based on the modified video swin transformer which can accept multi-source temporal remote sensing images as input and extract spatio-temporal features from various downsampling stages. Upon the pre-training of the foundation model, a versatile decoder is formulated. This decoder is capable of accepting features extracted from various stages of downsampling in the foundation model. It is specifically designed to generate a spectrum of crop mapping results, each correlating to its unique set of labels. The overall structure of the framework is shown in Figure 3.
@@ 2.3.1 Cropland Mapping, Field Boundary Delineation, and Early-season Crop Mapping
For cropland mapping and field boundary delineation tasks, we utilize the first 32 available scenes from the initial six months of each year as model inputs. In contrast, the early-season crop classification task employs a distinct temporal sampling strategy: we select 4 scenes per month to construct a 24-scene time series. This design serves dual purposes: (1) to explicitly validate our model’s capability in handling variable-length temporal inputs, and (2) to incorporate prior knowledge from previous year’s cropping patterns. The prior information, extracted using a ResNet50 (He et al., 2016) backbone, is concatenated with the foundation model’s output features. To maintain computational efficiency while ensuring feature richness, we deliberately limit the input sequence to 24 frames for this specific task. These experiments are designed to assess the effectiveness of our method in addressing high-resolution crop mapping tasks over extended temporal scales.
@@ 2.6 Implementation details
We develop our code using Python 3.11 and PyTorch-2.0. During the pre-training phase, the network learning rate is set to 1e-5. To ensure the stability of pre-training, the learning rate gradually increases from 1e-7 to 1e-5 over the first 5000 iterations, after which it gradually decreases throughout the pre-training process until it reaches 1e-6. The pre-training phase lasts for 50 epochs. The training is carried out on ten L40 GPUs, with a batch size of 80. The coefficient for the moving average is set to 0.001.
For specific crop mapping tasks, we employ the same network structure, that is, the unified crop mapping framework that we construct. The learning rate is set to 6e-5. All experiments are conducted on four L40 GPUs, with a batch size of 16, and last for 50 epochs. We select the model that performs best on the validation dataset as our final result.
@@ 4 Early-season Crop Mapping
Table 5: Performance comparison on early-season crop mapping (%). Metrics computed at field boundary pixels.
@@ 5 Paddy Rice Mapping in Monsoon Asia
Swin-based methods consistently outperform other architectures, with PIS achieving 86.49% F1 for rice detection and 90.32% average F1. This 3-4% advantage over CNN variants (best: CNN at 85.85% F1) and 4-5% over ViT models (best: SatMAE at 83.22% F1) confirms the superiority of hierarchical feature extraction for rice mapping. Notably, the randomly initialized Swin-scratch shows competitive performance (85.25% F1), suggesting the inherent suitability of Swin’s architecture for agricultural patterns.
@@
Considering the frequent cloud cover in the Monsoon Asia region, we incorporated MODIS data to aid in rice mapping. The experimental results suggest that the inclusion of MODIS data can enhance recall from 88.98% to 90.85%. This indicates that a larger proportion of rice can be correctly identified, emphasizing the effectiveness of our approach in extracting and integrating features from the MODIS input with those derived from HLS30. However, it can be observed that the precision decreases slightly with the inclusion of MODIS data. Although the F1 score still improves, the increase is not substantial. This is likely due to the significant difference in resolution between MODIS and HLS30. Furthermore, the accuracy of rice mapping using only HLS30 is already quite high, implying that in most scenarios, the best results can be achieved using only HLS30. Nevertheless, these findings still demonstrate that our proposed method can effectively integrate and process multi-source remote sensing data.
@@ 6 Winter Wheat Mapping in Asia
Interestingly, while winter wheat’s strong phenological characteristics make it relatively distinguishable from other crops, the performance gaps between methods are larger than those observed in rice mapping (Table 6). This accentuates our model’s specialized capabilities - the 8.8% F1 improvement over ViT models and 5.2% over CNNs demonstrate AgriFM’s particular aptitude for leveraging subtle temporal cues in coarse-resolution data.

### v2

@@ Abstract
Escalating climate change and global population growth demand precise crop mapping solutions to ensure food security. Accurate crop mapping fundamentally relies on modeling multi-scale spatiotemporal patterns, where spatial scales range from individual field textures to landscape-level context, and temporal scales capture both short-term phenological transitions and full growing-season dynamics, demanding robust joint spatiotemporal modeling capabilities. Existing approaches are inherently limited by their architectural designs that process spatial and temporal features separately, leading to significant information loss. Transformer-based remote sensing foundation models (RSFMs) offer promising potential for crop mapping due to their innate ability for unified spatiotemporal processing via self-attention mechanisms. However, current RSFMs remain suboptimal for crop mapping: they either employ fixed spatiotemporal windows that ignore the multi-scale nature of crop systems or completely disregard temporal information by focusing solely on spatial patterns. To bridge these gaps, we present AgriFM, a multi-source temporal remote sensing foundation model specifically designed for agricultural crop mapping. Our approach begins by establishing the necessity of simultaneous hierarchical spatiotemporal feature extraction, leading to the development of a modified Video Swin Transformer architecture where temporal down-sampling is synchronized with spatial down-sampling operations. This modified backbone enables efficient unified processing of long time-series satellite inputs while preserving critical multi-scale spatial patterns and phenological dynamics. AgriFM leverages temporally rich data streams from three satellite sources including MODIS, Landsat-8/9 and Sentinel-2, and is pre-trained on a global representative dataset comprising over 25 million image samples supervised by land cover products. The resulting framework incorporates a versatile decoder architecture that dynamically fuses these learned spatiotemporal representations, supporting diverse downstream tasks including cropland mapping, field-boundary delineation, early-season crop mapping, and specific crop mapping (e.g., winter wheat and paddy rice) with difference data sources. Comprehensive evaluations demonstrate AgriFM’s superior performance over conventional deep learning approaches and state-of-the-art general-purpose RSFMs across all downstream tasks. Notably, AgriFM demonstrates remarkable extensibility by supporting multi-source remote sensing data and variable temporal windows for diverse mapping tasks. Codes will be available at https://github.com/flyakon/AgriFM and https://glass.hku.hk.
@@ 1 Introduction
The growing demands for accurate crop mapping have spurred significant methodological evolution, beginning with phenology-based approaches. They utilize the distinct physical attributes of various crops, including their differing reflectance across spectral bands and their unique phenological traits at various growth phases. Examining and detecting the alterations throughout the crop growth cycle can efficiently differentiate various crops (Qiu et al., 2022; Dong et al., 2020; Qiu et al., 2017; Liang et al., 2024; Ashourloo et al., 2022). Qiu et al. (Qiu et al., 2022) introduced an innovative approach for generating yearly 500-m MODIS-derived national maps of various cropping systems in China, employing phenology-based mapping algorithms and pixel purity thresholds, resulting in an overall accuracy of 89%. Ashourloo et al. (Ashourloo et al., 2022) introduced a novel phenology-based approach utilizing Sentinel-2 time-series data for the efficient differentiation of wheat and barley across extensive regions, attaining an overall accuracy exceeding 76%. The issue with these methods is that they require setting different thresholds based on the mapping task and study area, which significantly affects their applicability. For complex scenarios and tasks, this may lead to erroneous or even failed mapping results.
@@
Transformer-based RSFMs have demonstrated superior capabilities in extracting spatiotemporal features across various remote sensing tasks, significantly outperforming traditional machine learning methods and deep learning architectures. This suggests RSFMs hold substantial potential for crop mapping, as they can theoretically model complete crop growth cycles end-to-end while integrating multi-source remote sensing data to improve mapping accuracy. Several pioneering studies have begun exploring RSFMs for crop mapping tasks. Fang et al. (Fang et al., 2025) developed a rice mapping method using NASA-IBM’s Prithvi foundation model (Jakubik et al., 2023) with Harmonized Landsat and Sentinel-2 (HLS30) data (Claverie et al., 2018) and successfully generated rice distribution maps across Monsoon Asia. Similarly, Qi et al. (Qin et al., 2025) proposed a spatiotemporal masking strategy for pretraining a spatiotemporal collaborative learning network (STCLN) to extract informative representations for crop mapping.
@@
While these approaches demonstrate promising results, they predominantly rely on Vision Transformer (ViT) architectures. This preference stems from ViT’s straightforward temporal data processing and relatively lower computational demands during masked image modeling pretraining. However, ViT’s inherent downsampling mechanism and fixed spatiotemporal patch windows may adversely affect crop mapping performance. Previous research (Chen et al., 2017; Huang et al., 2020) has proven pixel-wise classification tasks require high spatial feature fidelity, yet ViT’s patch embedding operation typically employs aggressive and fixed downsampling (e.g., 16×16 patches), potentially compromising subtle spatial-temporal differences. Although Fang and Qi’s implementations incorporate specialized modifications to mitigate these limitations, such adaptations increase methodological complexity while limiting generalizability.
@@
In contrast, Swin Transformer’s hierarchical feature extraction capability - resembling CNN’s multi-scale processing - appears better suited for crop mapping tasks. Like U-Net architectures, Swin Transformer can effectively fuse features across scales to enhance spatial precision (Cao et al., 2022). However, existing Swin-based RSFMs remain scarce and typically ignore temporal information, severely limiting their applicability for crop monitoring. While Video Swin Transformer (Liu et al., 2022) inherently supports joint spatiotemporal processing, few existing RSFMs have yet adopted it as the backbone architecture.
@@
To address these gaps, this paper first establishes the necessity of simultaneous hierarchical spatiotemporal feature extraction for crop mapping, leading to the development of a modified Video Swin Transformer (Liu et al., 2022) architecture where temporal down-sampling is synchronized with spatial down-sampling operations. This modified backbone enables efficient unified processing of long time-series satellite inputs while preserving critical multi-scale spatial patterns and phenological dynamics. We then develop AgriFM, a multi-source temporal remote sensing foundation model specifically designed for agricultural crop mapping.
@@
AgriFM leverages temporally rich data streams from MODIS, Landsat-8 / 9, and Sentinel-2 satellites, pre-trained on a global representative dataset with supervision from global land cover products. The pretraining dataset comprises over 25 million globally sampled images from MODIS (250m & 500m), Landsat-8/9 (30m), and Sentinel-2 (10m & 20m). For each satellite source, we randomly selected 16 chronological frames per year to capture complete seasonal variations. Unlike RSFMs relying on MIM or CL for pretraining, we incorporate land cover fraction from GLC_FCS30D data (Zhang et al., 2024c) for supervised pretraining. Specifically, we extract image-level land cover fractions within each sample area as regression targets, optimizing with L1 loss. This approach follows established practices demonstrating land cover priors’ effectiveness (Li et al., 2022; Wang et al., 2024). To address inherent label noise in reference products, we implemented an auxiliary teacher network with exponential moving average updates, effectively guiding the model toward robust feature learning while filtering out unreliable supervision signals.
@@
AgriFM incorporates a versatile decoder to dynamically fuse extracted multi-scale spatiotemporal representations, supporting diverse crop mapping tasks, including cropland mapping, field-boundary delineation, early-season crop mapping and specific crop mapping (e.g., winter wheat and paddy rice). For systematic evaluation, we compare AgriFM with three representative approaches: (1) ViT-based temporal RSFMs (Prithvi (Jakubik et al., 2023), SatMAE (Cong et al., 2022)), (2) Swin-based models without temporal processing (PIS (An et al., 2024)), and (3) conventional deep learning methods (CNNs / LSTMs). Experimental results demonstrate AgriFM’s consistent superiority, particularly in two crucial aspects: preserving fine-grained spatial details essential for field-level mapping, and effectively modeling long-term temporal patterns critical for crop type analysis - capabilities where existing methods show notable limitations. AgriFM achieves unified processing of heterogeneous data sources, variable temporal sequences, and diverse mapping tasks within a single network architecture. This comprehensive method overcomes the traditional need for task-specific model adaptations while delivering robust performance across various crop mapping tasks.
@@ 2.1.1 Multi-source Satellite Data for Pre-training
The MODIS data from the Terra satellite include the surface reflectance products MOD09A1 and MOD09Q1 with the temporal resolution of 8 days and the spatial resolutions of 500 m and 250 m, respectively. The MOD09Q1 product comprises the first two bands at 250m, and we extract bands 3-7 at 500m from the MOD09A1 product and then re-sample them to 250 meters. The MODIS data selected for pretraining are from 2020 and 2021.
@@ 2.1.3 Pre-training Dataset
After the pre-processing, we have constructed a multi-source temporal remote sensing pre-training dataset, which is derived from MODIS, Landsat 8/9, and Sentinel-2 satellite data at 250m&500m, 30m, and 10m&20m spatial resolution. To meet the input requirements of deep learning models, we randomly select these satellite data and crop the original satellite data into images of 224 x 224 pixels. For each image, a sequence of time series data is also generated with a minimum length of 16. Land cover fractions are computed to serve as pre-training supervision and the image with only background land cover is abandoned.
@@ 2.2 Method
In this section, we introduce our proposed multi-source temporal remote sensing foundation model AgriFM for crop mapping. AgriFM is built based on the modified video swin transformer which can accept multi-source temporal remote sensing images as input and extract spatio-temporal features from various downsampling stages. Upon the pre-training of the foundation model, a versatile decoder is formulated. This decoder is capable of accepting features extracted from various stages of downsampling in the foundation model. It is specifically designed to generate a spectrum of crop mapping results, each correlating to its unique set of labels. The overall structure of the framework is shown in Figure 3.
@@ 2.3.1 Cropland Mapping, Field Boundary Delineation, and Early-season Crop Mapping
For cropland mapping and field boundary delineation tasks, we utilize the first 32 available scenes from the first six months of each year as model inputs. In contrast, the early-season crop classification task employs a distinct temporal sampling strategy: we select 4 scenes per month to construct a 24-scene time series. This design serves dual purposes: (1) to explicitly validate our model’s capability in handling variable-length temporal inputs, and (2) to incorporate prior knowledge from previous year’s cropping patterns. The prior information, extracted using a ResNet50 (He et al., 2016) backbone, is concatenated with the foundation model’s output features. To maintain computational efficiency while ensuring feature richness, we deliberately limit the input sequence to 24 frames for this specific task. These experiments are designed to assess the effectiveness of our method in addressing high-resolution crop mapping tasks over extended temporal scales.
@@ 2.6 Implementation details
We develop our codes using Python 3.11 and PyTorch-2.0. During the pre-training phase, the network learning rate is set to 1e-5. To ensure the stability of pre-training, the learning rate gradually increases from 1e-7 to 1e-5 over the first 5000 iterations, after which it gradually decreases throughout the pre-training process until it reaches 1e-6. The pre-training phase lasts for 50 epochs. The training is carried out on ten L40 GPUs, with a batch size of 80. The coefficient for the moving average is set to 0.001.
For downstream crop mapping tasks, we employ the same network structure, that is, the unified crop mapping framework that we construct. The learning rate is set to 6e-5. All experiments are conducted on four L40 GPUs, with a batch size of 16, and last for 50 epochs. We select the model that performs best on the validation dataset as our final result.
@@ 4 Early-season Crop Mapping
Table 5: Performance comparison on early-season crop mapping (%).
@@ 5 Paddy Rice Mapping in Monsoon Asia
Swin-based methods consistently outperform other architectures, with PIS achieving 86.49% F1 and 90.32% average F1. This 3-4% advantage over CNN variants (best: CNN at 85.85% F1) and 4-5% over ViT models (best: SatMAE at 83.22% F1) confirms the superiority of hierarchical feature extraction for rice mapping. Notably, the randomly initialized Swin-scratch shows competitive performance (85.25% F1), suggesting the inherent suitability of Swin’s architecture for agricultural patterns.
@@
Considering the frequent cloud cover in the Monsoon Asia region, we incorporate MODIS data to aid in rice mapping. The experimental results suggest that the inclusion of MODIS data can enhance recall from 88.98% to 90.85%. This indicates that a larger proportion of rice can be correctly identified, emphasizing the effectiveness of our approach in extracting and integrating features from the MODIS input with those derived from HLS30. However, it can be observed that the precision decreases slightly with the inclusion of MODIS data. Although the F1 score still improves, the increase is not substantial. This is likely due to the significant difference in resolution between MODIS and HLS30. Furthermore, the accuracy of rice mapping using only HLS30 is already quite high, implying that in most scenarios, the best results can be achieved using only HLS30. Nevertheless, these findings still demonstrate that our proposed method can effectively integrate and process multi-source remote sensing data.
@@ 6 Winter Wheat Mapping in Asia
In addition, while winter wheat’s strong phenological characteristics make it relatively distinguishable from other crops, the performance gaps between methods are larger than those observed in rice mapping (Table 6). This accentuates our model’s specialized capabilities - the 8.8% F1 improvement over ViT models and 5.2% over CNNs demonstrate AgriFM’s particular aptitude for leveraging subtle temporal cues in coarse-resolution data.